# NRR-Core: Non-Resolution Reasoning as a Computational Framework for Contextual Identity and Ambiguity Preservation

### v7

@@ Abstract
Current artificial intelligence systems, despite remarkable capabilities in text generation and pattern recognition, exhibit a fundamental architectural limitation: they resolve ambiguity prematurely. This premature semantic collapse—the tendency to collapse multiple valid interpretations into a single output—stems from classical identity assumptions embedded in standard neural architectures. We propose Non-Resolution Reasoning (NRR), a computational framework that treats ambiguity retention as a valid reasoning mode rather than a defect to be eliminated. NRR introduces three core principles: (1) Non-Identity ($A\neq A$)—the same symbol refers to different entities across contexts; (2) Approximate Identity ($A\approx A$)—entities share partial structural overlap without being identical; and (3) Non-Resolution—conflicting interpretations can coexist without forced convergence. We formalize these principles through three architectural components: Multi-Vector Embeddings for context-dependent representation, Non-Collapsing Attention for parallel interpretation retention, and Contextual Identity Tracking (CIT) for maintaining $A\neq A$ across inference. We illustrate NRR’s potential through case studies in paradox handling, creative generation, and context-dependent reasoning. Crucially, we provide functional verification in a synthetic two-turn disambiguation task through a Turn 1 Entropy measurement demonstrating ambiguity preservation directly: NRR-lite maintains high entropy ($H=0.91$ bits, near-maximum $1.0$) at the ambiguous turn, while standard architectures collapse early ($H=0.15$ bits), demonstrating that NRR preserves interpretive flexibility until context arrives. NRR challenges the assumption that meaning must collapse to be useful, offering a foundation for AI systems capable of sophisticated ambiguity handling and creative reasoning. The question is not whether AI should resolve ambiguity, but when, how, and under whose control.
@@ 1.2 Contributions of This Work
    Theoretical Formalization: We provide a mathematical outline, including a formal state space and eight operators for non-collapsing computation, with explicit formalization of the $A\neq A\approx A$ principle.
@@ 5.5.1 State Space Formalization
$H(\mathcal{S})=-\sum_{i=1}^{n}p(i\mid\mathcal{S})\log p(i\mid\mathcal{S})$ (6)
@@ 5.5.2 Interference Operators
We introduce eight candidate interference operators that act on $\mathcal{S}$. We define an NRR interference operator as a transformation $T$ on $\mathcal{S}$ whose intended property is approximate information preservation, i.e., our design goal is that
@@
for small $\epsilon$. Table 1 specifies the operators and their roles; verifying and refining their information-preservation guarantees is left to future empirical and theoretical work. One operator, CPP Integration ($\kappa$), implements the Contradiction-Preservation Principle: contradictory hypotheses are retained rather than eliminated.
@@
Stripping $\sigma$ Removes learned biases inducing collapse
@@ 6.3 Results
Figure 2: Turn 1 Entropy Verification Results. Left: Entropy at Turn 1 (no context). NRR-lite maintains high entropy ($H\approx 0.91$ bits), near the theoretical maximum, while Baseline collapses ($H\approx 0.15$ bits). Center: Distribution of $P(\text{FINANCIAL})$ at Turn 1. NRR-lite centers around 0.5 (ambiguous); Baseline clusters near 0 or 1 (committed). Right: Both models achieve 100% accuracy when Turn 2 context is provided, confirming that NRR-lite’s ambiguity preservation does not impair resolution capability.
@@
    Robustness: Results replicate across 5 random seeds (paired $t$-test: $t(4)=13.3$, $p<0.001$).
@@ Acknowledgments
The author gratefully acknowledges the support of large language models—including ChatGPT (OpenAI), Claude (Anthropic), and Gemini (Google)—for their assistance in linguistic refinement, LaTeX formatting, and general proofreading support during the preparation of this manuscript. All concepts, arguments, and conclusions remain solely the responsibility of the author.

### v8

@@ Abstract
Current artificial intelligence systems, despite remarkable capabilities in text generation and pattern recognition, exhibit a fundamental architectural limitation: they resolve ambiguity prematurely. This premature semantic collapse—the tendency to collapse multiple valid interpretations into a single output—stems from classical identity assumptions embedded in standard neural architectures. We propose Non-Resolution Reasoning (NRR), a computational framework that treats ambiguity retention as a valid reasoning mode rather than a defect to be eliminated. NRR introduces three core principles: (1) Non-Identity ($A\neq A$)—the same symbol refers to different entities across contexts; (2) Approximate Identity ($A\approx A$)—entities share partial structural overlap without being identical; and (3) Non-Resolution—conflicting interpretations can coexist without forced convergence. We formalize these principles through three architectural components: Multi-Vector Embeddings for context-dependent representation, Non-Collapsing Attention for parallel interpretation retention, and Contextual Identity Tracking (CIT) for maintaining $A\neq A$ across inference. We provide functional verification of the Multi-Vector Embeddings component through a controlled synthetic task. We illustrate NRR’s potential through case studies in paradox handling, creative generation, and context-dependent reasoning. Crucially, we provide functional verification in a synthetic two-turn disambiguation task through a Turn 1 Entropy measurement demonstrating ambiguity preservation directly: NRR-lite maintains high entropy ($H=0.91$ bits, near-maximum $1.0$) at the ambiguous turn, while standard architectures collapse early ($H=0.15$ bits), demonstrating that NRR preserves interpretive flexibility until context arrives. NRR challenges the assumption that meaning must collapse to be useful, offering a foundation for AI systems capable of sophisticated ambiguity handling and creative reasoning. The question is not whether AI should resolve ambiguity, but when, how, and under whose control.
Implementation: A reference implementation accompanying this work is available at https://github.com/kei-saito-research/nrr-core.
@@ 1 Introduction
This paper is one module in an ongoing Non-Resolution Reasoning (NRR) research program, not a closed endpoint. Companion papers cover text-to-state mapping, implementation strategy, and cross-domain validation with distinct scopes. A series-level Program Map and update status will be maintained in repository documentation.
@@ 1.2 Contributions of This Work
    Theoretical Formalization: We provide a mathematical outline, including a formal state space, a compact core operator set for non-collapsing transitions, and extensible metadata functions for context tracking, with explicit formalization of the $A\neq A\approx A$ principle.
@@ 5.5.1 State Space Formalization
$H(\mathcal{S})=-\sum_{i=1}^{n}p(i\mid\mathcal{S})\log_{2}p(i\mid\mathcal{S})$ (6)
@@ 5.5.2 Interference Operators
We introduce a candidate family of interference operators that act on $\mathcal{S}$. The set is extensible rather than fixed; practical implementations may use a minimal subset. We define an NRR interference operator as a transformation $T$ on $\mathcal{S}$ whose intended property is approximate information preservation, i.e., our design goal is that
@@
for small $\epsilon$. Table 1 specifies the operators and their roles; verifying and refining their information-preservation guarantees is left to future empirical and theoretical work. Notation note: Later implementation papers distinguish state-level and item-level operators. In particular, IME/Universal use item-level updates (often denoted $\sigma_{\text{item}},\delta_{\text{item}}$), while this table lists the broader state-level candidate family. One operator, CPP Integration ($\kappa$), implements the Contradiction-Preservation Principle: contradictory hypotheses are retained rather than eliminated.
@@
State-level Calibration $\sigma$ Uniform scaling/check operator (may be entropy-invariant)
@@ 6.3 Results
Figure 2: Turn 1 Entropy Verification Results (aggregated across 5 seeds). Left: Entropy at Turn 1 (no context). NRR-lite maintains high entropy ($H\approx 0.91$ bits), near the theoretical maximum, while Baseline collapses ($H\approx 0.15$ bits). Center: Distribution of $P(\text{FINANCIAL})$ at Turn 1 across all test samples (500 total: 5 seeds × 100 samples each). NRR-lite centers around 0.5 (ambiguous); Baseline clusters near 1.0 (committed to financial interpretation). Right: Per-seed Turn 1 Entropy breakdown across all 5 random seeds, demonstrating consistent preservation by NRR-lite (blue, $H\approx 0.9$ bits across seeds) and variable collapse by Baseline (red, ranging from 0.01 to 0.37 bits depending on initialization).
@@
    Robustness: Results replicate across 5 random seeds (paired $t$-test: $t(4)=11.9$, $p<0.001$).
  6. 6.
    Baseline Instability: Baseline’s high standard deviation ($\text{SD}=0.13$, 87% of mean) reflects architectural instability across random initializations, with entropy ranging from 0.01 to 0.37 bits depending on seed. This contrasts sharply with NRR-lite’s consistent performance ($\text{SD}=0.04$, 4% of mean), demonstrating robust ambiguity preservation independent of initialization.
@@ Acknowledgments
The author gratefully acknowledges the support of large language models—including Claude (Anthropic), ChatGPT and Codex (OpenAI), and Gemini (Google)—for their assistance in linguistic refinement, LaTeX formatting, and general proofreading support during the preparation of this manuscript. All concepts, arguments, and conclusions remain solely the responsibility of the author.